OpenStudy (anonymous):

This is an absolute value equation I am a bit confused with 4 I 3a+4 I = 16a+8 between 4 and 3 and again between 4 and= and the absolute value signs.

OpenStudy (anonymous):

the absolute signs means the absolute value of 3a + 4 - that means the positive value only so that if the value of 3a + 4 is negative it takes on the positive value. so for example if a = -10 making 3a+4 = -26 - its abso;ute value is positive 26

OpenStudy (anonymous):

4 I 3a+4 I = 16a+8 is the complete equation

OpenStudy (anonymous):

4 I 3a+4 I = 16a+8 divide both sides by 4 I 3a+4 I = 4a + 2 take positive value : 3a + 4 = 4a + 2 a = 2 is one solution take negative value: 3a + 4 = -(4a + 2) 3a + 4 = -4a - 2 7a = -6 a = -6/7 is other solution

OpenStudy (anonymous):

4 I 3a+4 I = 16a+8 divide both sides by 4 I 3a+4 I = 4a + 2 take positive value : 3a + 4 = 4a + 2...........I subtracted 4 from both sides to Isolate the left 3a so I end up with 4a -2 then after dividing both sides by 4a i arrive at a = -2 a = 2 is one solution

OpenStudy (anonymous):

The second negative - answer I came up as you did.